The one-sample t statistic for a test of H0: μ = 11 vs. Ha: μ < 11 based on n = 13 observations has the test statistic value of t = −1.25. What is the p-value for this test?

a) 0.418

b) 0.882

c) 0.000

d) 0.118

e) 0.235
Answer #1

this is the left tailed test .  

The null and alternative hypothesis is ,

H0 :   = 11

Ha : <

Test statistic = t =−1.25 df=n-1=13-1=12

P(t ,df < -1.25,12) =  0.1176

P-value = 0.118